# Capillary levelling of thin liquid films of power-law rheology

### 21 Feb 2024 v1

@@ Abstract
We find solutions that describe the levelling of a thin fluid film coating a substrate, comprising a non-Newtonian power law fluid, under the influence of surface tension. We consider the evolution from both periodic and localised initial conditions. These solutions exhibit the generic property that the profiles are weakly singular (that is, higher order derivatives do not exist), at points where the pressure gradient vanishes. Numerical solutions of the thin film equation are shown to approach one of the two cases for appropriate initial conditions.
@@ 1 Introduction
The importance of free-surface lubrication or coating flows to industry and science is well documented [Craster2009, Oron1997, Weinstein2004]. Examples of applications include the construction of smooth, or deliberately patterned, surfaces, describing the evolution of lava flows and ice sheets, among others. The mathematical modelling of such phenomena using the lubrication approximation, which allows the full governing partial differential equations and boundary conditions to be collapsed to a single equation for the film thickness, is also ubiquitous [Craster2009, Oron1997].
For small scale industrial applications in which surface tension is a dominant effect, fluids are frequently better modelled using a non-Newtonian, rather than Newtonian, rheological model. A very popular model is the power-law model:
@@
where $\tau$ is the stress tensor, $\dot{\gamma}$ is the strain rate tensor, $n$ is the power law rheology exponent, and $\alpha$ is a constant of proportionality (see [Myers2005] for a discussion of different rheological models). The Newtonian case is recovered when $n=1$, in which case $\alpha$ is the fluid viscosity. When $n<1$, the effective viscosity is decreasing in the strain rate, while for $n>1$, it is increasing. These two cases are thus known as shear-thinning and shear-thickening, respectively. Shear thinning fluids are more common in practice. The importance of rheology in the coating industry is described in [Eley2005].
From a mathematical perspective, the evolution of a lubricating film, comprising a power law fluid, under the effect of surface tension, has been considered with a focus on droplet spreading [King2001], or inclined or vertical substrates [Allouche2015, Balmforth2003, Sylvester1973]. The competition between surface tension and other effects such as van der Waals disjoining pressure, [Garg2017], and horizontally induced thermocapillary stress [Mantripragada2022].
In this article we focus on the deceptively simple problem that is the levelling of a film due to surface tension, in which a perturbed uniform film tends to flatten out over time. In the Newtonian case, this phenomenon has been studied in the context of viscometry [Benzaquen2013, Benzaquen2014, McGraw2011, McGraw2012]; by measuring the evolution of a perturbed film over time, the viscosity of the fluid may be determined. In these works, explicit similarity solutions are found that model the levelling process. This same application in the power-law fluid case has been studied to a lesser extent; [Iyer1996] perform numerical calculations of the levelling problem with a Carreau fluid and compare to experimental results. [Ahmed2015] perform a purely numerical study on power law fluids. The similarity solutions considered in the present study are also a case of nonlinear differential equations whose existence is established in [Bernis1991]. However, the explicit computation of similarity solutions that govern levelling behaviour, and comparison with numerical solutions, has not previously been undertaken in for power-law rheology.
@@
where we have explicitly defined the flux $q$. Here $n$ is the power law exponent in the rheology; $n<1$ $n=1$, and $n>1$ correspond to the shear thinning, Newtonian, and shear thickening cases, respectively. In the derivation of (2), the third derivative arises as the gradient of the Laplace pressure (the interface curvature being given by the second derivative in the lubrication approximation).
The governing equation (2) has as an exact solution a perfectly uniform film which can be taken to be $h=1$ identically (we are free in the nondimensionalisation to specify the characteristic thickness as the vertical length scale). In this work we consider the evolution of perturbations of this uniform film, distinguishing two cases: the evolution of a perturbation that is periodic in space, which is equivalent to a perturbation over an entire finite domain with no-flux boundary conditions; and the evolution of an initially localised perturbation in an infinite spatial domain (see Fig. 1). These two cases generalise the results of a Newtonian liquid film, which are modal linear stability of a flat interface, and the similarity solution derived in [Benzaquen2013, Benzaquen2014], respectively. We numerically determine the similarity solutions that describe each of these in the non-Newtonian case. The computed solutions demonstrate a general property of solutions to power-law thin film equations, namely, the non-existence of higher derivatives (depending on the exponent $n$). As well as of mathematical interest, this property is important to establish from a modelling perspective, as it affects the accuracy and potentially the correctness of numerical schemes. We also carry out numerical simulations of the fully time-dependent thin film equation, and show that solutions with more general initial conditions are attracted to the periodic or similarity solutions as appropriate.
@@ 2.1 Weakly singular nature
In the Newtonian case, (2) is degenerate in that the flux vanishes at points where the thickness $h$ vanishes. In the non-Newtonian case, (2) is degenerate even if $h$ is not zero, at a point $x_{0}$ where the pressure gradient (that is, the third derivative of $h$) is zero. A sensible expansion for $h$ near such a point is
$h=h_{0}(t)+h_{1}(t)(x-x_{0})+h_{2}(t)(x-x_{0})^{2}+h_{r}(t)|x-x_{0}|^{r},\qquad r\geq 3,$
where $r$ is an as-yet unknown power. On substitution into (2), the leading order term in the flux $q$ is
$q=h_{0}^{2+1/n}\left[r(r-1)(r-2)|h_{r}||x-x_{0}|^{r-3}\right]^{1/n}\operatorname{sgn}(x-x_{0})+\ldots$
@@
$\dot{h}_{0}+h_{0}^{2+1/n}\left[r(r-1)(r-2)|h_{r}|\right]^{1/n}\left(\frac{r-3}{n}\right)|x-x_{0}|^{(r-3)/n-1}+\ldots=0.$
Generally $\dot{h}_{0}=h_{t}(x_{0},t)\neq 0$ at such a point, thus, we thus require $r=3+n$. For non-integer $n$, we expect the film profile will be weakly singular, such that $h(x,t)$ is $3+\lfloor n\rfloor$ times differentiable. This is particular important for shear-thinning flows ($n<1$) as then we only expect weak, rather than classical, solutions of the fourth order equation (2) to exist. We will observe this singular behaviour in each of the solutions constructed in this article. We note that as special cases, if $\dot{h}_{0}=0$, then a singularity at higher order will occur as the gradient of flux must balance with another term, while if $n$ is an integer not equal to unity, then there will be higher order noninteger powers in the flux that need to be balanced, requiring higher order non-integer powers in the expansion of $h$. We do not consider these special cases further here.
@@ 2.3 Periodic initial condition
where the amplitude $\bar{A}$ is arbitrary (equivalent to time-translational invariance). For $n\neq 1$, the near-level behaviour is essentially nonlinear, and linear stability approaches are not appropriate. Instead, we assume a solution of (4) more generally in which the time and space dependence may be separated, or, equivalently, a similarity solution in which the spatial scale remains fixed. The appropriate ansatz depends on whether the fluid is shear thinning or thickening:
@@
where $\bar{H}(x)$ is a periodic, but as yet undetermined, function (the $n$-dependent constant is included to simplify the problem for $\bar{H}$). The above ansatz represents infinite-time but algebraic decay for the shear thinning ($n<1$) case, and finite-time (at $T=0$) levelling for the shear thickening $n>1$ case; of course, the solutions are invariant to translations in time, and the finite levelling time will depend on the initial amplitude.
@@
$[|\bar{H}^{\prime\prime\prime}|^{1/n}\operatorname{sgn}(\bar{H}^{\prime\prime\prime})]^{\prime}=\bar{H},$
for both shear-thinning and thickening. Let $U=|\bar{H}^{\prime\prime\prime}|^{1/n}\operatorname{sgn}(\bar{H}^{\prime\prime\prime})$, then $U$ satisfies
@@ 2.3.1 Numerical computation
Corresponding to the behaviour of the original problem near a singular point described in Section 2.1 for noninteger $n$, solutions to (6) will be weakly singular at points $x_{0}$ where $U(x_{0})=0$, with an expansion of the form
$U=U_{1}(x-x_{0})+U_{2}(x-x_{0})^{2}+U_{3}(x-x_{0})^{3}+U_{r}|x-x_{0}|^{4+n}\operatorname{sgn}(x-x_{0}).$
In the shear-thinning case, $U$ will thus be four (but not five)-times differentiable, corresponding to the profile $\bar{H}$ (and so the time-dependent profile $h$) being three but not four-times differentiable at points where the pressure gradient $h_{xxx}=0$. However, since (6) is fourth order, this singularity is weak enough to be able to proceed (that is, we have essentially integrated the equation by solving for the antiderivative of $\bar{H}$).
Assuming symmetry, we find the solution on the half period $0<x<\pi$ with the conditions
@@
In the half-period, $U$ may be taken to be positive. We calculate solutions to the boundary value problem (6), (7) numerically using the MATLAB function bvp4c. An initial guess is provided close to $n=1$ from the asymptotic solution described below (Section 2.3.2). A continuous branch of solutions is then constructed by numerical continuation for decreasing and increasing power law exponent $n$, using the previously computed solution as an initial guess for the next.
In Fig. 2 we plot the results of this computation, focussing on the values $n=1/2$ and $n=3/2$. We include both the profiles $\bar{H}=U^{\prime}$, as well as the fourth derivative $\bar{H}^{\prime\prime\prime\prime}=n|U|^{n-1}U^{\prime}$. The profiles themselves are visually difficult to distinguish from the sinusoidal profiles that are exact for the Newtonian case ($n=1$), but have non-arbitrary, $n$-dependent amplitude. In plotting the fourth derivative, the difference between the cases is stark, in particular, that the shear thinning case ($n=1/2$) is singular at the peak and trough of the periodic profile ($x=0,\pi$), while for the shear-thickening case ($n=3/2$), the fourth derivative is bounded and continuous (although the next derivative will indeed be singular). We also plot the amplitude $\bar{A}$, calculated as half of the peak to trough distance $\bar{A}=(\bar{H}(0)-\bar{H}(\pi))/2$; this value varies only weakly in the range of $n$ values calculated.
Figure 2: Spatial profiles $\bar{H}(x)$ representing levelling solutions for a periodic initial condition, $n=1/2$ (shear-thinning) and $n=3/2$ (shear-thickening). (a) The solution profiles (calculated from solutions $U$ to (6) with $\bar{H}=U^{\prime}$). These profiles are close to sinusoidal. (b) The fourth derivative $\bar{H}^{\prime\prime\prime\prime}$ show the singular nature of these profiles, with $\bar{H}$ only three times differentiable for $n=1/2$, and only four times differentiable for $n=3/2$. (c) The dependence of amplitude $\bar{A}$ on power law exponent $n$.
2.3.2 Close to Newtonian limit
@@
$U_{0}=\bar{A}\sin(x),$
with the amplitude $\bar{A}$ determined by a solvability condition for $U_{1}$. Applying the same boundary conditions to $U_{1}$, we must have that (by the Fredholm alternative)
@@ 2.4 Localised initial condition
subject to symmetry conditions at zero. This equation may be integrated once and the boundary conditions used to result in the third-order equation
@@
A third condition is imposed by requiring solutions to decay as $\eta\to\infty$. Similarly to the periodic case, since we only deal with third derivatives of $F$ in (10), the singularities that are present (in particular in the fourth derivative for $n<1$) will not prevent numerical computation of the solution.
In the Newtonian case, the similarity solution has been previously identified in [Benzaquen2013, Benzaquen2014]; In this case the similarity exponent $\alpha=1/4$, and the equation for the profile (10) is linear:
$F^{\prime\prime\prime}=\eta F$
which, given $F(0)=1$, $F^{\prime}(0)=0$, and decay at infinity, has exact solution in terms of hypergeometric functions:
$F(\eta)={}_{0}F_{2}\left([],[1/2,3/4],\eta^{4}/64\right)-\frac{\Gamma(3/4)}{4\Gamma(5/4)}\eta^{2}{}_{0}F_{2}\left([],[5/4,3/2],\eta^{4}/64\right).$
For $n\neq 1$, numerical solutions to (10) are obtained by shooting from $\eta=0$. Since $F(0)=1,F^{\prime}(0)=0$, there is one parameter $F^{\prime\prime}(0)$ that must be determined numerically by requiring that $F\to 0$ as $\eta\to\infty$. We use ode45 in MATLAB to solve the equation up to a sufficiently large value of $\eta$, combined with a root finding method to determine the appropriate value of $F^{\prime\prime}(0)$. The solutions so found are plotted in Fig. 3, including both the profiles $F$ and the fourth derivative, calculated from the numerical solution using
@@
As for periodic solutions, plotting this fourth derivative shows the presence of singularities at higher order. As the similarity solution oscillates an infinite number of times as $\eta\to\infty$, there are an infinite sequence of such singularities.
Figure 3: Similarity solutions for thin film levelling of an initially localised perturbation, for power law exponents $n=1/2$ and $n=3/2$. While the solution profiles (a) appear qualitatively similar, plotting the fourth derivative (b) demonstrates the weakly singular nature of the similarity solutions.

### 10 Apr 2024 v2

@@ Abstract
We find solutions that describe the levelling of a thin fluid film, comprising a non-Newtonian power-law fluid, that coats a substrate and evolves under the influence of surface tension. We consider the evolution from both periodic and localised initial conditions as separate cases. Particular (similarity) solutions in each of these two cases exhibit the generic property that the profiles are weakly singular (that is, higher-order derivatives do not exist), at points where the pressure gradient vanishes. Numerical simulations of the thin film equation, with either periodic or localised initial condition, are shown to approach the appropriate particular solution.
@@ 1 Introduction
The importance of free-surface lubrication or coating flows to industry and science is well documented [Craster2009, Oron1997, Weinstein2004]. Examples of applications include the construction of smooth (or deliberately patterned) surfaces, as well as predicting the evolution of lava flows and ice sheets, among others. The mathematical modelling of such phenomena using the lubrication approximation is also ubiquitous [Craster2009, Oron1997]; this approximation allows the full governing partial differential equations and interfacial boundary conditions to be collapsed to a single equation for the film thickness, with only boundary conditions in the horizontal coordinates needed.
For small-scale industrial applications in which surface tension is a dominant effect, fluids are frequently better modelled using a non-Newtonian, rather than Newtonian, rheological model. A very popular model is the power-law model:
@@
where $\tau$ is the stress tensor, $\dot{\gamma}$ is the strain rate tensor, $n$ is the power-law rheology exponent, and $\alpha$ is a constant of proportionality (see [Myers2005] for a discussion of different rheological models). The Newtonian case is recovered when $n=1$, in which case $\alpha$ is the fluid viscosity. When $n<1$, the effective viscosity is decreasing in the strain rate, while for $n>1$, it is increasing. These two cases are thus known as shear-thinning and shear-thickening, respectively. Shear-thinning fluids are more common in practice. The importance of rheology in the coating industry is described in [Eley2005].
From a mathematical perspective, the evolution of a lubricating film, comprising a power-law fluid, under the effect of surface tension, has been considered with a focus on droplet spreading [King2001, Rafai2004], and on inclined or vertical substrates [Allouche2015, Balmforth2003, Sylvester1973]. The competition between surface tension and other effects such as van der Waals disjoining pressure, [Garg2017], and horizontally induced thermocapillary stress [Mantripragada2022], have also been considered with the use of numerical simulations.
In this article we focus on the deceptively simple problem that is the levelling of a film due to surface tension, in which a perturbed uniform film tends to flatten out over time. In the Newtonian case, this phenomenon has been studied in the context of viscometry [Benzaquen2013, Benzaquen2014, McGraw2011, McGraw2012]; by measuring the evolution of a perturbed film over time, the viscosity of the fluid may be determined. In these works, explicit similarity solutions are found that model the levelling process. This same application in the power-law fluid case has been studied to a lesser extent; [Iyer1996] perform numerical calculations of the levelling problem with a Carreau fluid and compare to experimental results. [Ahmed2015] perform a purely numerical study on power-law fluids. However, the explicit computation of similarity solutions that govern levelling behaviour, and comparison with numerical solutions, has not previously been undertaken for the power-law rheology.
@@
where we have explicitly defined the flux $q$. Here $n$ is the power-law exponent in the rheology; $n<1$ $n=1$, and $n>1$ correspond to the shear-thinning, Newtonian, and shear-thickening cases, respectively. In the derivation of (2), the third derivative arises as the gradient of the Laplace pressure (the interface curvature being given by the second derivative in the lubrication approximation). The governing equation (2) has as an exact solution a perfectly uniform film which can be taken to be $h=1$ identically (we are free in the nondimensionalisation to specify the characteristic thickness as the vertical length scale).
In this work we consider the evolution of perturbations of a uniform film, distinguishing two cases: the evolution of a perturbation that is periodic in space, and the evolution of an initially localised perturbation in an infinite spatial domain (see Fig. 1). The periodic case is equivalent to a perturbation over a finite domain with width equal to the period. These two cases generalise known results for a Newtonian liquid film, namely, modal linear stability of a flat interface for periodic perturbations, and the similarity solution derived in [Benzaquen2013, Benzaquen2014] for localised perturbations. In the following section we will semi-analytically determine universal behaviour in each case as a perturbed film becomes level, before comparing with numerical simulations in Section 3.
@@ 2.1 Weakly singular nature
In the Newtonian case, the thin film equation (2) is only singular at points where the thickness $h$ vanishes, which are not present in the study of levelling. In the non-Newtonian case, (2) is singular even if $h$ is not zero, at a point $x_{0}$ where the pressure gradient (that is, the third derivative of $h$) is zero. We demonstrate this singular nature by expanding a solution near such a point. A sensible expansion for $h$ near $x_{0}$ is
$h=h_{0}(t)+h_{1}(t)(x-x_{0})+h_{2}(t)(x-x_{0})^{2}+h_{r}(t)|x-x_{0}|^{r}+\ldots,\qquad r\geq 3,$
where $r$ is an as-yet unknown power. On substitution into (2), the leading-order term in the flux $q$ is
$q=h_{0}^{2+1/n}\left[r(r-1)(r-2)|h_{r}||x-x_{0}|^{r-3}\right]^{1/n}\operatorname{sgn}(h_{r})\operatorname{sgn}(x-x_{0})+\ldots$
@@
$\dot{h}_{0}+h_{0}^{2+1/n}\left[r(r-1)(r-2)|h_{r}|\right]^{1/n}\left(\frac{r-3}{n}\right)\operatorname{sgn}(h_{r})|x-x_{0}|^{(r-3)/n-1}+\ldots=0.$
Here $\dot{h}_{0}$ is the time derivative of the leading coefficient function. Generally $\dot{h}_{0}\neq 0$ at such a point, which requires $r=3+n$. For non-integer $n$, we expect the film profile will be weakly singular, in the sense that $h(x,t)$ is only $3+\lfloor n\rfloor$ times differentiable. This is particularly important for shear-thinning flows ($n<1$) given the equation (2) is fourth order. We will observe this singular behaviour in each of the solutions constructed in this article. We note that as special cases, if $\dot{h}_{0}=0$, then a singularity at higher order will occur as the gradient of flux must balance with another term, while if $n$ is an integer not equal to unity, then there will be higher-order noninteger powers in the flux that need to be balanced, requiring higher-order non-integer powers in the expansion of $h$. We do not consider these special cases further here.
@@ 2.3 Periodic initial condition
where the amplitude $\bar{A}$ is arbitrary (equivalent to time-translational invariance). For $n\neq 1$, the near-level behaviour is essentially nonlinear, and linear stability approaches are not appropriate. Instead, we assume a solution of (4) more generally in which the time and space dependence may be separated, or, equivalently, a similarity solution in which the horizontal spatial scale remains fixed. The appropriate ansatz depends on whether the fluid is shear-thinning or thickening:
@@
where $\bar{H}(x)$ is a periodic, but as yet undetermined, function (the $n$-dependent constant is included to simplify the problem for $\bar{H}$). The above ansatz represents infinite-time but algebraic decay as $T\to\infty$ for the shear-thinning ($n<1$) case, and finite-time (at $T\to 0^{-}$) levelling for the shear-thickening $n>1$ case. As written (5) is valid for $T>0$ and $T<0$ for $n<1$ and $n>1$, respectively; however, the solutions are of course invariant to translations in time, and in particular the finite levelling time will depend on the initial amplitude.
@@
$[|\bar{H}^{\prime\prime\prime}|^{1/n}\operatorname{sgn}(\bar{H}^{\prime\prime\prime})]^{\prime}=\bar{H}$ (6)
(where ${}^{\prime}$ represents differentiation with respect to $x$), for both shear-thinning and thickening. Let $U=|\bar{H}^{\prime\prime\prime}|^{1/n}\operatorname{sgn}(\bar{H}^{\prime\prime\prime})$, then on differentiating (6) three times, $U$ satisfies
@@ 2.3.1 Numerical computation
Solutions to (7) will be weakly singular at points $x_{0}$ where $U(x_{0})=0$, with an expansion of the form
$U=C_{1}(x-x_{0})+C_{2}(x-x_{0})^{2}+C_{3}(x-x_{0})^{3}+C_{r}|x-x_{0}|^{4+n}\operatorname{sgn}(x-x_{0})+\ldots,$
where $C_{1},C_{2},\ldots$, are the coefficients in the series. This singularity corresponds to the behaviour of the original problem near a singular point described in Section 2.1 for noninteger $n$. In particular, in the shear-thinning case, $U$ will be four (but not five)-times differentiable, corresponding to the profile $\bar{H}$ (and so the time-dependent profile $h$) being three but not four-times differentiable at points where the pressure gradient $h_{xxx}=0$. However, since (7) is fourth order, this singularity is weak enough to be able to proceed (that is, we have essentially integrated the equation by solving for the antiderivative of $\bar{H}$).
Assuming a solution profile that is symmetric about a point $x=0$, a $2\pi$-periodic solution is equivalent to finding a solution on the half period $0<x<\pi$ that satisfies the conditions
@@
In the half-period, $U$ may be taken to be positive. We calculate solutions to the boundary value problem (7), (8) numerically using the MATLAB function bvp4c. An initial guess is provided close to $n=1$ from the asymptotic solution described below (Section 2.3.2). A continuous branch of solutions is then constructed by numerical continuation for decreasing and increasing power-law exponent $n$, using the previously computed solution as an initial guess for the next.
In Fig. 2 we plot the results of this computation, focussing on the values $n=1/2$ and $n=3/2$. These exponents are chosen as the examples of shear-thinning and thickening cases throughout our study. We include both the profiles $\bar{H}=U^{\prime}$, as well as the fourth derivative $\bar{H}^{\prime\prime\prime\prime}=n|U|^{n-1}U^{\prime}$. The profiles themselves are visually difficult to distinguish from the sinusoidal profiles that are exact for the Newtonian case ($n=1$), but have non-arbitrary, $n$-dependent amplitude. In plotting the fourth derivative, the difference between the cases is stark, in particular, that the shear-thinning case ($n=1/2$) is singular at the peak and trough of the periodic profile ($x=0,\pi$), while for the shear-thickening case ($n=3/2$), the fourth derivative is bounded and continuous (although the next derivative will indeed be singular). We also plot the amplitude $\bar{A}$, calculated as half of the peak to trough distance $\bar{A}=(\bar{H}(0)-\bar{H}(\pi))/2$; this value varies only weakly in the range of $n$ values calculated.
Figure 2: Spatial profiles $\bar{H}(x)$ representing levelling solutions for a periodic initial condition, $n=1/2$ (shear-thinning) and $n=3/2$ (shear-thickening), with the exact cosine profile of the Newtonian case $n=1$ shown for reference. (a) The solution profiles (calculated from solutions $U$ to (7) with $\bar{H}=U^{\prime}$). These profiles are close to sinusoidal. (b) The fourth derivative $\bar{H}^{\prime\prime\prime\prime}$ show the singular nature of these profiles, with $\bar{H}$ only three times differentiable for $n=1/2$, and only four times differentiable for $n=3/2$. (c) The dependence of amplitude $\bar{A}$ on power-law exponent $n$.
2.3.2 Close-to-Newtonian limit
@@
$U_{0}=\bar{A}\sin(x).$
The amplitude $\bar{A}$ is determined from a solvability condition for $U_{1}$. Applying the same boundary conditions to $U_{1}$, we must have that (by the Fredholm alternative)
@@ 2.4 Localised initial condition
(here ${}^{\prime}$ refers to derivatives with respect to $\eta$) subject to symmetry conditions at zero. This equation may be integrated once and the boundary conditions used to result in the third-order equation
@@
A third condition is imposed by requiring solutions to decay as $\eta\to\infty$. We note that (11) is a special case of a class of nonlinear differential equations whose existence is established in [Bernis1991], although they do not explicitly construct solutions numerically. Similarly to the periodic case, since we only deal with third derivatives of $F$ in (11), the singularities that are present (in particular in the fourth derivative for $n<1$) will not prevent numerical computation of the solution.
In the Newtonian case, the similarity solution has been previously identified in [Benzaquen2013, Benzaquen2014]; the similarity exponent is $\alpha=1/4$, and the equation for the profile (11) is linear:
$F^{\prime\prime\prime}=\eta F.$ (12)
Given $F(0)=1$, $F^{\prime}(0)=0$, and decay at infinity, (12) has an exact solution that may be expressed in terms of the Fourier integral
$F(\eta)=C\int_{-\infty}^{\infty}\exp\left(\mathrm{i}\xi\eta-\frac{\xi^{4}}{4}\right)\,\mathrm{d}\xi=C\eta^{1/3}\int_{-\infty}^{\infty}\exp\left[\left(\mathrm{i}\zeta-\frac{\zeta^{4}}{4}\right)\eta^{4/3}\right]\,\mathrm{d}\zeta,$ (13)
where $C=\sqrt{2}/\Gamma(1/4)$ is determined from the condition $F(0)=1$. The large-$\eta$ asymptotic behaviour of (13) is readily determined from a steepest descent calculation, with a contour that passes through two critical points $\zeta=\mathrm{e}^{\mathrm{i}\pi/6},\mathrm{e}^{5\mathrm{i}\pi/6}$, which ultimately results in
$F(\eta)\sim\frac{4\sqrt{\pi}}{\sqrt{3}\Gamma(1/4)}\eta^{-1/3}\exp\left(-\frac{3}{8}\eta^{4/3}\right)\cos\left(\frac{3\sqrt{3}}{8}\eta^{4/3}-\frac{\pi}{6}\right),\qquad\eta\to\infty.$
This asymptotic result indicates that $F$ decays in oscillatory fashion as $\eta\to\infty$. In addition, the solution (13) may be represented exactly in terms of hypergeometric functions:
$F(\eta)={}_{0}F_{2}\left([],[1/2,3/4],\eta^{4}/64\right)-\frac{\Gamma(3/4)}{4\Gamma(5/4)}\eta^{2}{}_{0}F_{2}\left([],[5/4,3/2],\eta^{4}/64\right).$ (14)
This solution is depicted in Figure 3a.
Returning to the non-Newtonian case $n\neq 1$, numerical solutions to (11) are obtained by shooting from $\eta=0$. Since $F(0)=1,F^{\prime}(0)=0$, there is one parameter $F^{\prime\prime}(0)$ that must be determined numerically by requiring that $F\to 0$ as $\eta\to\infty$. We use ode45 in MATLAB to solve the equation up to a sufficiently large value of $\eta$, combined with a root finding method to determine the appropriate value of $F^{\prime\prime}(0)$. We perform this calculation for the exponents $n=1/2$ and $n=3/2$ (for $n=1$, the numerical method matches the exact solution (14)). The solutions so found are plotted in Fig. 3, including both the profiles $F$ and the fourth derivative, calculated from the numerical solution using
@@
As was the case for periodic solutions, plotting this fourth derivative shows the presence of singularities at higher order. Similar to the Newtonian case, the non-Newtonian similarity solutions oscillate an infinite number of times as $\eta\to\infty$, so there is an infinite sequence of such singularities.
Figure 3: Similarity solutions for thin film levelling of an initially localised perturbation, for power-law exponents $n=1/2$ and $n=3/2$, as well as the Newtonian case $n=1$. While the solution profiles (a) appear qualitatively similar, plotting the fourth derivative (b) demonstrates singularities in the higher derivatives of the non-Newtonian similarity solutions.